OpenStudy (anonymous):

Find the values of the trigonometric functions of θ from the information given. sin θ = 12/13 θ in Quadrant II

OpenStudy (anonymous):

Use the 5-12-13 right triangle and sign conventions in Quadrant II (sine is positive while cosine and tangent are negative). \[ \cos \theta = - \frac {5}{13} \]\[ \tan \theta = - \frac {12}{5} \]\[ \csc \theta = \frac {13}{12} \]\[ \sec \theta = - \frac {13}{5} \]\[ \cot \theta = - \frac {5}{12} \]

OpenStudy (anonymous):

.